Assume that • U.S. population is 328 million people


Assume that 

• U.S. population is 328 million people 

    • 70 million individuals are legally classified as unable to work (are less than 16 years of age, institutionalized, or in the military) 

    • 96 million who could be working are classified as uninterested working 

    • 7 million are classified as unemployed 


Given this information, what is the unemployment rate? 

State your answer as a percentage (type in the percentage symbol), and round to one decimal place, i.e. x.x% 
Answer #1

Unemployment rate is 4.3%...

Total population of U.S = 328 million..

Total Labour force

= Total population of U.S - (legally unable to work + uninterested to work)

= 328 -(70+96) = 328 - 166 = 162 million

unemployed people= 7 million

Hence, Unemployment rate

= (Unemployed people /Labour force)*100

= (7/162)*100 = 4.3%...